# MMEarth: Exploring Multi-Modal Pretext Tasks For Geospatial Representation Learning

### v1

@@ Abstract
The volume of unlabelled Earth observation (EO) data is huge, but many important applications lack labelled training data. However, EO data offers the unique opportunity to pair data from different modalities and sensors automatically based on geographic location and time, at virtually no human labor cost. We seize this opportunity to create a diverse multi-modal pretraining dataset at global scale. Using this new corpus of 1.2 million locations, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach to learn general-purpose representations for optical satellite images. Our approach builds on the ConvNeXt V2 architecture, a fully convolutional masked autoencoder (MAE). Drawing upon a suite of multi-modal pretext tasks, we demonstrate that our MP-MAE approach outperforms both MAEs pretrained on ImageNet and MAEs pretrained on domain-specific satellite images. This is shown on several downstream tasks including image classification and semantic segmentation. We find that multi-modal pretraining notably improves the linear probing performance, e.g. 4pp on BigEarthNet and 16pp on So2Sat, compared to pretraining on optical satellite images only. We show that this also leads to better label and parameter efficiency which are crucial aspects in global scale applications.  [^1] The MMEarth dataset is available on the project page: vishalned.github.io/mmearth. The dataset construction code is available here: github.com/vishalned/MMEarth-data. The MP-MAE code for training and evaluation is available here: github.com/vishalned/MMEarth-train.
@@ 1 Introduction
Learning representations through self-supervision has seen major advancements in recent years [5, 22, 33, 32, 2]. This gives hope that even applications with limited supervision data may profit from the promises of the deep learning revolution. However, some self-supervised techniques still fall short when evaluated on image domains other than ImageNet, e.g., ground-level images from the natural world [47]. Thus, to develop generalizing representation learning techniques, it is necessary to have suitable large-scale datasets of different domains. As an alternative to “pure” self-supervision paradigm, multi-modal data offer great potential for learning good semantic representations [36, 4, 31]. Aligned multi-modal datasets are key for advancing two major research directions in computer vision: i) exploiting multi-modal data for representation learning, and ii) advancing representation learning to exploit multi-modal data for inference.
@@
One domain that is particularly well-suited for studying the potential of multi-modal data is Earth Observation (EO). Geolocated data from different sensors can automatically be aligned at virtually no human labor cost. Aligning such EO data at large scale became substantially easier thanks to cloud-based platforms like Google Earth Engine [21] and Microsoft Planetary Computer [30]. Furthermore, real-world remote sensing applications are often hampered by a lack of high-quality reference data. Human annotation of satellite data is not always possible from the images alone, for example when mapping carbon stocks or species abundance. Such applications require field measurements by experts that are costly, time consuming, and do not scale to large areas. Other high-quality measurement systems, such as airborne or spaceborne LiDAR, can provide a means to scale reference data, but still provide a biased sample of the Earth, due to the limits of global sampling. Here, we present a global dataset called MMEarth at the scale of ImageNet-1k [13] with multi-modal data for 1.2M locations. We also release our data collection framework that allows to collect data from different sensors given a list of geospatial coordinates. While MMEarth can be used to advance both aforementioned research directions, we demonstrate its potential for the first goal of exploiting multi-modal data for improving representation learning. Specifically, our goal is to learn general-purpose representations for optical satellite images from the Sentinel-2 mission that are predictive for a wide range of downstream tasks including crop type, landcover, and climate zone classification. Thus, in contrast to other work in this area, we neither aim to generalize to other sensor inputs such as Scale-MAE [37] nor to make use of multiple modalities at inference during downstream tasks as MultiMAE [4] or 4M [31]. As shown by previous work, specializing a model to a particular input modality can be beneficial compared to models that aim to generalize to multiple input modalities [31].
@@
To explore the potential of multiple pretext tasks, we build on fully convolutional masked autoencoders (FCMAE) and extend the ConvNeXt V2-MAE approach [52] with multi-modal reconstruction tasks for pretraining. As opposed to MAEs based on vision transformers, the FCMAE is well suited for downstream tasks that require different input image sizes, which is a inevitable when evaluating the representations on existing EO benchmarks [27]. Intuitively, such a multi-task pretraining strategy with a shared encoder should lead to representations that generalize better to new downstream tasks that are not known at pretraining time. We evaluate the potential for new downstream tasks on the GEO-Bench [27] Sentinel-2 tasks including image-level multi-class as well as multi-label classification and semantic segmentation. Our contributions can be summarized as follows:
@@
    Method - We explore the effect of multi-modal pretext tasks for both pixel-level and image-level modalities when learning general purpose representations for interpreting optical images from Sentinel-2. Therefore, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach to use MMEarth to learn better representations for Sentinel-2 optical images.
@@
    Results - We show that pretraining MP-MAE on MMEarth outperforms not only MAEs pretrained on ImageNet but also on Sentinel-2 images. While both fine-tuning and linear probing performance is improved, linear probing profits substantially from the multi-modal pretext tasks.
@@ 2.1 Masked image modelling
Masked Image Modelling (MIM) can see its roots in denoising autoencoders that are trained to reconstruct original images from partially corrupted inputs [49]. This idea was extended to pretrain CNN autoencoders by reconstructing a large area in an image with the aim to learn better representations by conditioning a portion of the image on its surroundings [34]. With the success of masked language modelling (for example, BERT [14]) and the rise of Vision Transformers (ViT) [16], Masked Autoencoder (MAE) approaches have been developed that reconstruct masked input patches [16, 22, 53]. When masking a large fraction of the input sequence, MIM is not only a successful pretext task to learn semantic representations, but also leads to a more efficient ViT encoder [22, 53]. This masking strategy has also been extended to spatio-temporal modelling to learn from videos [18, 44, 51, 6] as well as to learn representations for 3D point clouds [55]. To combine this random masking strategy with the inductive bias of CNNs, we use ConvNeXt V2 [52] which is a Fully Convolutional MAE (FCMAE). The encoder is implemented with sparse convolutions [9] to improve efficiency like in the ViT-based MAE. ConvNeXt V2 also uses a learnable mask token which serves as the input to the convolutional decoder. Instead of using self-attention, the decoder is based on dense convolutions with kernels that cover the full spatial extent of the image. This architecture design pretrained on ImageNet, leads to a high-performing model family of different sizes, ranging from 3.7M (‘Atto’) to 650M parameters (‘Huge’). We adapt the ConvNeXt V2 architecture design to satellite images and pretrain on our MMEarth dataset to demonstrated the potential of multi-modal EO data for representation learning.
@@ 2.2 Multi-modal representation learning
Multi sensory information has been argued to help humans (and animals) to understand the natural world [40], which motivates the development of machine learning approaches that learn representations from multi-modal data. Successful approaches such as CLIP [36] are based on a contrastive learning objective while learning separate encoders for each input modality. This approach has been extended to learn from images, text, and audio extracted from videos [1]. While contrastive approaches usually rely on aligned samples from various modalities, non-contrastive approaches were proposed to e.g. learn image-text representations with a single encoder to avoid the need for paired image-text data [19]. It has been shown that visual tasks are (more or less) related and that it can be more data efficient to learn multiple tasks jointly [56]. In line with these observations, multi-task self-training [20] leverages pseudo-labels for multi-task learning to yield improved general purpose representations. In addition to such multi-task objectives, MIM has been extended to the multi-modal image setting [51, 4, 31]. Masked feature prediction [51] has been demonstrated as an effective pretext task, e.g. by reconstructing the Histograms of Oriented Gradients (HOG) instead of the original image, arguing that the local contrast normalization plays a crucial role. MultiMAE [4] extends the MAE approach [22] to multiple input and output modalities using RGB, semantic label masks, and depth. Recently, 4M [31] proposed a masked modelling framework to train a single autoencoder for a range of input and output modalities. Instead of reconstructing each modality in the original space (e.g. pixels for RGB), each modality is first mapped to discrete tokens using a modality-specific autoencoder. These tokens are then used as the input and output to a transformer-based autoencoder trained with masked modelling leading to impressive any-to-any reconstruction results. However, their ablation study demonstrates that pretraining with a single input modality can outperform a model trained with multiple input modalities. While previous works have focused on image modalities that provide pixel-level data, our approach makes use of both pixel-level and image-level modalities.
@@ 2.3 Representation learning in Earth observation
In applications where labelled training data is abundant, supervised learning has been used in a range of applications like tree segmentation [46], marine debris detection [39], or snow depth estimation [12]. However, only few studies have scaled supervised deep learning to global scales to map land cover [8], canopy height [28], or solar panels [26]. Additionally, large annotated EO datasets have been created to replace ImageNet with a domain-specific pretraining dataset[10, 41, 42, 50, 7]. Some bi-modal datasets provide Sentinel-2 optical and Sentinel-1 SAR images in Europe [42] and globally around cities [50]. In contrast, MMEarth provides a global pretraining dataset with 12 modalities balanced across 14 biomes [15].
@@
Geolocation. In an approach called geography-aware self-supervised learning (GASSL) [3], the contrastive learning objective leverages the spatio-temporal structure of EO data. Using spatially aligned images, positive image pairs can be created by using observations at different times, avoiding the need of artificial augmentations. In addition, localization is used as a pretext task by predicting the clusters of sampled locations. SatCLIP [25] focuses on learning location embeddings in a contrastive manner by learning a Sentinel-2 encoder and a location encoder with a contrastive objective. Time. Seasonal contrast (SeCo [29]) is another temporal contrastive approach that learns multiple embedding subspaces that are invariant to either seasonal changes, synthetic augmentations, or all transformations. Alternatively, image time series can be exploited in masked modelling like demonstrated in SatMAE [11]. Presto [45] is a masked modelling approach that focuses on learning representations for multi-modal time series using a light-weight transformer architecture, which only relies on pixel time series inputs, but ignores textural features. Modality. Scale-MAE [37] aims to generalize to different input resolutions by conditioning on the ground sampling distance (GSD, size of a pixel on the ground). Most related to our work is SatlasPretrain [7], a large multi-task dataset for optical satellite images including Sentinel-2. Their goal was “to label everything that is visible in a satellite image” using various data sources (Open Street Map, lidar scans, landcover maps) and human annotators (domain experts and Mechanical Turk workers). Our MMEarth dataset creation differs, as it is fully automatic without relying on human annotators and also considers modalities that are not explicitly visible for humans in the optical images. Such modalities include everything that cannot be annotated by a human like radar, temperature, precipitation, or canopy height. We hypothesize that predicting such cross-modal relations from images requires the extraction of semantic features.
All these works provide important insights in how to make better use of satellite data and its metadata in representation learning. Our work extends along the last aspect — exploiting multiple modalities — whereas space and time serve as the underlying structure to pair data from different sources automatically without relying on human annotation. Furthermore, we treat geolocation and time as two additional modalities in our framework.
@@ 3 MMEarth dataset
Table 1: Modalities in the MMEarth dataset. Each location provides 12 aligned modalities with a total of 46 bands. The upper half of the modalities provides dense data at the pixel-level and the bottom half at the image-level.
@@
MMEarth contains data for 1.2 million locations distributed around the world, making its optical image count comparable to ImageNet-1K [13]. At each location, data from 12 geo-aligned modalities were collected, grouped into pixel-level and image-level modalities, see Table 1. The six pixel-level modalities represent raster data of size $128\times 128$ pixels which capture 1.28 km×$\times$1.28 km on the ground (e.g., Sentinel-2, Sentinel-1, Aster DEM, Dynamic World, and ESA World Cover). The remaining six image-level modalities represent scalar values for each location (e.g., Biome, Ecoregion, ERA5 temperature, ERA5 precipitation, Geolocation, and Sentinel-2 observation date). The size of the input images is large enough for typical downstream EO tasks  [27] and yields a reasonable dataset volume. MMEarth was sampled uniformly across 14 biomes (Fig. 1, e.g., Mangroves, Temperate Conifer Forests, Tundra, etc.) [15]. To increase diversity, we considered data from the four years 2017–2020 (Fig. 1). Further, we ensured that time-critical modalities were collected around the Sentinel-2 observation date, which serves as the reference. Additionally, all pixel-level modalities were re-projected to the Sentinel-2 10m grid, if needed, yielding a harmonized data cube. In addition to the full dataset, we provide two ‘taster’ subsets to facilitate research in multi-modal representation learning with limited compute resources: MMEarth100k (100k locations at 128$\times$128 pixels) and MMEarth64 (1.2M locations center cropped to 64$\times$64 pixels), see supplementary material for details. All the data were downloaded from Google Earth Engine (GEE) [21]. In the following subsections, we provide a brief overview of the pixel-level and image-level modalities listed in Table 1. More details in the supplementary material.
@@ 3.1 Pixel-level modalities
Sentinel-2. An optical satellite mission that provides global coverage at least every 5 days at 10m resolution. Its multi-spectral sensor comprises of 12 bands, with four bands RGBN (RGB and near infrared) at 10m and near and short-wave infrared bands at 20m and 60m resolution, respectively. We upsample all bands to the 10m grid. Sentinel-2 provides two processing levels: top of atmosphere reflectance (L1C) and the bottom of atmosphere reflectance (L2A). L2A is atmospherically corrected, harmonizing the spectral data, but is not yet available for all years on a global scale. To ensure global coverage for the years 2017–2020, MMEarth includes both L2A and L1C data. Sentinel-1. Synthetic Aperture Radar (SAR) data at 10m resolution, collected in two viewing directions along an ascending and descending orbit, resulting in four bands each. Aster GDEM. A Global Digital Elevation Model derived from the Advanced Spaceborne Thermal Emission and Reflection Radiometer (ASTER). The Aster GDEM data covers 99% of the Earth’s landmass at 15m resolution. We include elevation and slope. ETH-GCHM. A 10m Global Canopy Top Height Map [28] for the year 2020 derived from Sentinel-2 and sparse vertical structure data from GEDI [17], a spaceborne lidar mission. Dynamic World. A land cover dataset that contains semantic segmentation of Sentinel-2 images for nine categories at 10m resolution. We aggregate the per image maps for each year to reduce noise in the data. ESA World Cover. A 10m global land cover map based on Sentinel-1 and Sentinel-2 data with 11 categories for the year 2020.

### v2

@@ Abstract
The volume of unlabelled Earth observation (EO) data is huge, but many important applications lack labelled training data. However, EO data offers the unique opportunity to pair data from different modalities and sensors automatically based on geographic location and time, at virtually no human labor cost. We seize this opportunity to create MMEarth, a diverse multi-modal pretraining dataset at global scale. Using this new corpus of 1.2 million locations, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach to learn general-purpose representations for optical satellite images. Our approach builds on the ConvNeXt V2 architecture, a fully convolutional masked autoencoder (MAE). Drawing upon a suite of multi-modal pretext tasks, we demonstrate that our MP-MAE approach outperforms both MAEs pretrained on ImageNet and MAEs pretrained on domain-specific satellite images. This is shown on several downstream tasks including image classification and semantic segmentation. We find that pretraining with multi-modal pretext tasks notably improves the linear probing performance compared to pretraining on optical satellite images only. This also leads to better label efficiency and parameter efficiency which are crucial aspects in global scale applications.  [^1] The MMEarth dataset is available on the project page: vishalned.github.io/mmearth. The dataset construction code is available here: github.com/vishalned/MMEarth-data. The MP-MAE code for training and evaluation is available here: github.com/vishalned/MMEarth-train.
@@ 1 Introduction
Learning representations through self-supervision has seen major advancements in recent years [5, 22, 34, 32, 2]. This gives hope that even applications with limited supervision data may profit from the promises of the deep learning revolution. However, some self-supervised techniques still fall short when evaluated on image domains other than ImageNet, e.g., ground-level images from the natural world [48]. Thus, to develop generalizing representation learning techniques, it is necessary to have suitable large-scale datasets from different domains. As an alternative to the “pure” self-supervision paradigm, multi-modal data offer great potential for learning good semantic representations [37, 4, 31]. Aligned multi-modal datasets are key for advancing two major research directions in computer vision: i) exploiting multi-modal data for representation learning, and ii) advancing representation learning to exploit multi-modal data for inference.
@@
One domain that is particularly well-suited for studying the potential of multi-modal data is Earth Observation (EO). Geolocated data from different sensors can automatically be aligned at virtually no human labor cost. Aligning such EO data at large scale has become substantially easier thanks to cloud-based platforms like Google Earth Engine [21] and Microsoft Planetary Computer [30]. Furthermore, real-world remote sensing applications are often hampered by a lack of high-quality reference data. Human annotation of satellite data is not always possible from the images alone, for example when mapping carbon stocks or species abundance. Such applications require field measurements by experts that are costly, time consuming, and do not scale to large areas. Other high-quality measurement systems, such as airborne or spaceborne LiDAR, can provide a means to scale reference data, but still provide a biased sample of the Earth, due to the limits of global sampling. Here, we present a global dataset called MMEarth at the scale of ImageNet-1k [13] with multi-modal data for 1.2M locations. We also release our data collection framework for collecting data from different sensors given a list of geospatial coordinates. While MMEarth can be used to advance both aforementioned research directions, we demonstrate its potential for the first goal of exploiting multi-modal data for improving representation learning. Specifically, our goal is to learn general-purpose representations for optical satellite images from the Sentinel-2 mission that are predictive for a wide range of downstream tasks including crop type, landcover, and climate zone classification. Thus, in contrast to other work in this area, we neither aim to generalize to other sensor inputs such as Scale-MAE [38] nor to make use of multiple modalities during inference in downstream tasks such as MultiMAE [4] or 4M [31]. As shown by previous work, specializing a model to a particular input modality can be beneficial compared to models that aim to generalize to multiple input modalities [31].
@@
To explore the potential of multiple pretext tasks, we build on fully convolutional masked autoencoders (FCMAE) and extend the ConvNeXt V2-MAE approach [53] with multi-modal reconstruction tasks for pretraining. As opposed to MAEs based on vision transformers, the FCMAE is well suited for downstream tasks that require different input image sizes, which is inevitable when evaluating the representations on existing EO benchmarks [27]. Intuitively, such a multi-task pretraining strategy with a shared encoder should lead to representations that generalize better to new downstream tasks that are not known at pretraining time. We evaluate the generalization potential using Sentinel-2 downstream tasks from GEO-Bench [27], including image-level multi-class and multi-label classification as well as pixel-level semantic segmentation tasks. Our contributions can be summarized as follows:
@@
    Method - We explore the effect of multi-modal pretext tasks for both pixel-level and image-level modalities when learning general purpose representations for interpreting optical images from Sentinel-2. Therefore, we propose a Multi-Pretext Masked Autoencoder (MP-MAE) approach to using MMEarth to learn better representations for Sentinel-2 optical images.
@@
    Results - We show that pretraining MP-MAE on MMEarth outperforms not only MAEs pretrained on ImageNet but also on Sentinel-2 images. While both fine-tuning and linear probing performance are improved, linear probing benefits particularly from the multi-modal pretext tasks.
@@ 2.1 Masked image modelling
Masked Image Modelling (MIM) can see its roots in denoising autoencoders that are trained to reconstruct original images from partially corrupted inputs [50]. This idea was extended to pretrain CNN autoencoders by reconstructing a large area in an image with the aim of learning better representations by conditioning a portion of the image on its surroundings [35]. With the success of masked language modelling (for example, BERT [14]) and the rise of Vision Transformers (ViT) [16], Masked Autoencoder (MAE) approaches have been developed that reconstruct masked input patches [16, 22, 54]. When masking a large fraction of the input sequence, MIM is not only a successful pretext task to learn semantic representations, but also leads to a more efficient ViT encoder [22, 54]. This masking strategy has also been extended to spatio-temporal modelling to learn from videos [18, 45, 52, 6] as well as to learn representations for 3D point clouds [56]. To combine this random masking strategy with the inductive bias of CNNs, we use ConvNeXt V2 [53], which is a Fully Convolutional MAE (FCMAE). The encoder is implemented with sparse convolutions [9] to improve efficiency like in the ViT-based MAE. ConvNeXt V2 also uses a learnable mask token which serves as the input to the convolutional decoder. Instead of using self-attention, the decoder is based on dense convolutions with kernels that cover the full spatial extent of the image. This architecture design pretrained on ImageNet yields a high-performing model family of different sizes, ranging from 3.7M (‘Atto’) to 650M parameters (‘Huge’). We adapt the ConvNeXt V2 architecture design to satellite images and pretrain on our MMEarth dataset to demonstrate the potential of multi-modal EO data for representation learning.
@@ 2.2 Multi-modal representation learning
Multi-sensory information has been argued to help humans (and animals) to understand the natural world [41], which motivates the development of machine learning approaches that learn representations from multi-modal data. Successful approaches such as CLIP [37] are based on a contrastive learning objective while learning separate encoders for each input modality. This approach has been extended to learn from images, text, and audio extracted from videos [1]. While contrastive approaches usually rely on aligned samples from various modalities, non-contrastive approaches were proposed to e.g. learn image-text representations with a single encoder to avoid the need for paired image-text data [19]. It has been shown that visual tasks are (more or less) related and that it can be more data-efficient to learn multiple tasks jointly [57]. In line with these observations, multi-task self-training [20] leverages pseudo-labels for multi-task learning to yield improved general purpose representations. In addition to such multi-task objectives, MIM has been extended to the multi-modal image setting [52, 4, 31]. Masked feature prediction [52] has been demonstrated as an effective pretext task, e.g. by reconstructing the Histograms of Oriented Gradients (HOG) instead of the original image, arguing that the local contrast normalization plays a crucial role. MultiMAE [4] extends the MAE approach [22] to multiple input and output modalities using RGB, semantic label masks, and depth. Recently, 4M [31] proposed a masked modelling framework to train a single autoencoder for a range of input and output modalities. Instead of reconstructing each modality in the original space (e.g. pixels for RGB), each modality is first mapped to discrete tokens using a modality-specific autoencoder. These tokens are then used as the input and output to a transformer-based autoencoder trained with masked modelling, leading to impressive any-to-any reconstruction results. However, their ablation study demonstrates that pretraining with a single input modality can outperform a model trained with multiple input modalities. While previous works have focused on image modalities that provide pixel-level data, our approach makes use of both pixel-level and image-level modalities.
@@ 2.3 Representation learning in Earth observation
In applications where labelled training data is abundant, supervised learning has been used in a range of applications like tree segmentation [47], marine debris detection [40], and snow depth estimation [12]. However, only a few studies have scaled supervised deep learning to global scales to map land cover [8], canopy height [28], or solar panels [26]. Additionally, large annotated EO datasets have been created to replace ImageNet with a domain-specific pretraining dataset[10, 42, 43, 51, 7]. Some bi-modal datasets provide Sentinel-2 optical and Sentinel-1 SAR images in Europe [43] and globally around cities [51]. The benefit of multi-modal input data has been studied for landcover mapping at the scale of Europe in a fully supervised setting [33]. In contrast, MMEarth provides a global pretraining dataset with 12 modalities balanced across 14 biomes [15].
@@
Geolocation. In an approach called geography-aware self-supervised learning (GASSL) [3], the contrastive learning objective leverages the spatio-temporal structure of EO data. Using spatially aligned images, positive image pairs can be created by using observations at different times, avoiding the need for artificial augmentations. In addition, localization is used as a pretext task by predicting the clusters of sampled locations. SatCLIP [25] focuses on learning location embeddings in a contrastive manner by learning a Sentinel-2 encoder and a location encoder with a contrastive objective. Time. Seasonal contrast (SeCo [29]) is another temporal contrastive approach that learns multiple embedding subspaces that are invariant to either seasonal changes, synthetic augmentations, or all transformations. Alternatively, image time series can be exploited in masked modelling as was demonstrated in SatMAE [11]. Presto [46] is a masked modelling approach that focuses on learning representations for multi-modal time series using a light-weight transformer architecture that only relies on pixel time series inputs and ignores textural features. Modality. Scale-MAE [38] aims to generalize to different input resolutions by conditioning on the ground sampling distance (GSD, size of a pixel on the ground). Most related to our work is SatlasPretrain [7], a large multi-task dataset for optical satellite images including Sentinel-2. Their goal was “to label everything that is visible in a satellite image” using various data sources (Open Street Map, lidar scans, landcover maps) and human annotators (domain experts and Mechanical Turk workers). Our MMEarth dataset creation differs, as it is fully automatic without relying on human annotators and also considers modalities that are not explicitly visible to humans in optical images. Such modalities include everything that cannot be annotated by a human like radar, temperature, precipitation, or canopy height. We hypothesize that predicting such cross-modal relations from images requires the extraction of semantic features.
All these works provide important insights into how to make better use of satellite data and its metadata in representation learning. Our work extends along the last aspect — exploiting multiple modalities — whereas space and time serve as the underlying structure for pairing data from different sources automatically without relying on human annotation. Furthermore, we treat geolocation and time as two additional modalities in our framework.
@@ 3 MMEarth dataset
Table 1: Modalities in the MMEarth dataset. Each location provides 12 aligned modalities with a total of 46 bands. The upper half of the modalities provide dense data at the pixel level and the bottom half at the image level.
@@
MMEarth contains data for 1.2 million locations distributed around the world, making its optical image count comparable to that of ImageNet-1K [13]. At each location, data from 12 geo-aligned modalities were collected, grouped into pixel-level and image-level modalities as per Table 1. The six pixel-level modalities represent raster data of size $128\times 128$ pixels which capture $1.28\times 1.28$ km on the ground (e.g., Sentinel-2, Sentinel-1, Aster DEM, Dynamic World, and ESA World Cover). The remaining six image-level modalities represent scalar values for each location (e.g., Biome, Ecoregion, ERA5 temperature, ERA5 precipitation, Geolocation, and Sentinel-2 observation date). The size of the input images is large enough for typical downstream EO tasks  [27] and yields a reasonable dataset volume. MMEarth was sampled uniformly across 14 biomes (Fig. 1, e.g., Mangroves, Temperate Conifer Forests, Tundra, etc.) [15]. To increase diversity, we considered data from the four years 2017–2020 (Fig. 1). Further, we ensured that time-critical modalities were collected around the Sentinel-2 observation date, which serves as the reference. Additionally, all pixel-level modalities were re-projected to the Sentinel-2 10m grid, if needed, yielding a harmonized data cube. In addition to the full dataset, we provide two subsets to facilitate research in multi-modal representation learning with limited compute resources: MMEarth100k (100k locations at $128\times 128$ pixels) and MMEarth64 (1.2M locations center-cropped to $64\times 64$ pixels); see supplementary material for details. All the data were downloaded from Google Earth Engine (GEE) [21]. In the following subsections, we provide a brief overview of the pixel-level and image-level modalities listed in Table 1. More details are provided in the supplementary material.
@@ 3.1 Pixel-level modalities
Sentinel-2. An optical satellite mission that provides global coverage at least every 5 days at 10m resolution. Its multi-spectral sensor comprises of 12 bands, with four bands RGBN (RGB and near infrared) at 10m and near and short-wave infrared bands at 20m and 60m resolution, respectively. We upsample all bands to the 10m grid. Sentinel-2 provides two processing levels: top of atmosphere reflectance (L1C) and bottom of atmosphere reflectance (L2A). L2A is atmospherically corrected, harmonizing the spectral data, but is not yet available for all years on a global scale. To ensure global coverage for the years 2017–2020, MMEarth includes both L2A and L1C data. Sentinel-1. Synthetic Aperture Radar (SAR) data at 10m resolution, collected in two viewing directions along an ascending and descending orbit, resulting in four bands each. Aster GDEM. A Global Digital Elevation Model derived from the Advanced Spaceborne Thermal Emission and Reflection Radiometer (ASTER). The Aster GDEM data covers 99% of the Earth’s landmass at 15m resolution. We include elevation and slope. ETH-GCHM. A 10m Global Canopy Top Height Map [28] for the year 2020 derived from Sentinel-2 and sparse vertical structure data from GEDI [17], a spaceborne LiDAR mission. Dynamic World. A land cover dataset that contains semantic segmentation of Sentinel-2 images for nine categories at 10m resolution. We aggregate the per-image maps for each year to reduce noise in the data. ESA World Cover. A 10m global land cover map based on Sentinel-1 and Sentinel-2 data with 11 categories for the year 2020.